Molly's sister is three years younger than twice her age. If the sum of their ages is 24, what is each of their ages?
Answer by stanbon(75887)

Molly's sister is three years younger than twice her age. If the sum of their ages is 24, what is each of their ages?
-----------------
Equations:
s = 2m - 3
s + m = 24
------
Substitute for "s" and solve for "m"::
2m-3 + m = 24
3m = 27
m = 9 (Molly's age now)
------
Solve for "s":
s = 2m -3
s = 18-3 = 15 yrs (sister's age now)
------------
Cheers,
Stan H.
----------